OpenStudy (anonymous):

find a) area and b) circumference of a circle that has a diameter of 40 cm. use 3.14 for pi.

OpenStudy (anonymous):

C≈125.66cm

OpenStudy (anonymous):

circumference is 125.66

OpenStudy (anonymous):

is that answer you were looking for?

OpenStudy (anonymous):

thank you. Can you please explain how you got that(:

OpenStudy (anonymous):

@imxbatman

OpenStudy (anonymous):

\[c = \pi \times d\]

OpenStudy (anonymous):

so \[c = 3.14 \times 40\]

OpenStudy (anonymous):

thank you sure @ChiefArnav

OpenStudy (anonymous):

sir*

OpenStudy (anonymous):

movin go its \[c = 31.4 \times 4\]

OpenStudy (anonymous):

therefore \[c = 125.6 cm\]

OpenStudy (anonymous):

for the area there are two formulae which do you want us to use @mayita_01?

OpenStudy (anonymous):

as \[d \times d\]

OpenStudy (anonymous):

d is 40 so 40 times 40 is 1600

OpenStudy (anonymous):

so now it's \[1/4 \times 3.14 \times 1600\]

OpenStudy (anonymous):

so its \[3.14 \times 400\]

OpenStudy (anonymous):

as \[1/4 \times 1600 = 400\]

OpenStudy (anonymous):

so now its \[314 \times 4\]

OpenStudy (anonymous):

so its 314 times 1 = 314 314 times 2 = 628 314 times 3 = 942 314 times 4 = 1256 therefore the area is 1256cm